OpenStudy (anonymous):

can somebody do this: log 6x + 3 =2 ,then solve for x

OpenStudy (anonymous):

Isolate the log 6x term

OpenStudy (anonymous):

Subtract 3 both side to do that

OpenStudy (anonymous):

log 6x=-1

OpenStudy (anonymous):

Now change it into exponent form

OpenStudy (anonymous):

remember that the 3 is contained to 6x +3

OpenStudy (anonymous):

and it is attached to log of base 10

OpenStudy (anonymous):

how do you comment on that

OpenStudy (anonymous):

log 6x+3=-1 change into exponent form 10^-1=6x+3 .1=6x+3

OpenStudy (anonymous):

so, i think the question is log(6x+3) = 2; this is equivalent to: log(6x+3) = log (100) 6x+3 = 100 6x = 97 x = 97/6 www.oojih.com